OpenStudy (immanuelv):

Every month, a girl gets allowance. Assume last year she had no money, and kept it up to now. Then she spends 1/2 of her money on clothes, then 1/3 of the remaining money on games, and then 1/4 of the remaining money on toys. After she bought all of that, she had $7777 left. Assuming she only gets money by allowance, how much money does she earn every month?

OpenStudy (anonymous):

she's loaded

OpenStudy (anonymous):

i'ts only may and she had over 14k damn that girl is spoiled

OpenStudy (dumbcow):

Let x be the monthly allowance Let M be money spent 12x -M = 7777 M = (1/2)(12x) + (1/3)(1/2)(12x) + (1/4)(2/3)(1/2)(12x) M = 6x +2x+x M = 9x 12x -9x = 7777 3x = 7777 x = 2592.33